# Hierarchical Empowerment: Towards Tractable Empowerment-Based Skill-Learning

### v1

@@ 4.2 Results
We present our results in a few ways. Figure 3 provides charts for the phase two performances for most of the experiments. Results of the three versus four level Ant experiment is provided in section A of the Appendix. We also provide a video presentation showing sample episodes of the trained agents at the following url: https://www.youtube.com/watch?v=kwLUYMsMRNI. In addition, we provide before and after images of the learned goal spaces in section C of the Appendix.
The experimental results support both contributions of the framework. Goal-Conditioned Empowerment and DIAYN were compared in the Ant Field – Small and Half Cheetah – Small domains. Per the phase two results in 3, Goal-Conditioned Empowerment was able to complete both tasks while DIAYN was not able to solve either task. As we describe section F of the Appendix, in our implementation of DIAYN we tracked the scale of the learned variational distribution $q_{\phi}(z|s_{0},s_{t})$. As we hypothesized, the standard deviation of the variational distribution did not decrease over time, indicating that the the skills were not specializing and targeting specific states. On the other hand, per the performance charts and video, Goal-Conditioned Empowerment was able to learn skills to achieve goals in the phase two goal space.
The experimental results also support the need for hierarchy to compute the long-term empowerment of states. In both the two versus three level agents (i.e., one skill level vs. two skill level agents) and in the three versus four level comparison, the agents with the additional level outperformed, often significantly, the agent with fewer levels. In addition, the surface area of the skills learned by the four level Hierarchical Empowerment agents is significantly larger than has been reported in prior empowerment-based skill-learning papers. In our largest Ant domain solved by the four level agent, the phase 2 goal space with size 800x800 is over four orders of magnitude larger than the surface area reported in the DIAYN Ant task (4x4), and over 2 orders of magnitude larger than the area (30x30) partially achieved by Dynamics-Aware Unsupervised Discovery of Skills (DADS) [Sharma et al., 2020]. Section B of the Appendix provides a visual of the differences in surface area coverage among Hierarchical Empowerment, DADS, and DIAYN. In terms of wall clock time, the longest goals in the 800x800 domain required over 13 minutes of movement by a reasonably efficient policy.
@@ 5 Related Work
There is an extensive history of skill-learning approaches [Parr and Russell, 1997, Sutton et al., 1999, Dietterich, 1999, Stolle and Precup, 2002, Simsek et al., 2005, Konidaris and Barto, 2009, Bacon et al., 2017]. Here we review the categories of skill-learning most closely related to our proposed framework.
Empowerment. There are several works that take a different approach to maximizing the mutual information of the skill channel. DADS [Sharma et al., 2020] optimizes the symmetric version of the mutual information of the skill channel: $H(S_{n}|s_{0})-H(S_{n}|s_{0},Z)$. Similar to the empowerment-based skill-learning methods mentioned previously, DADS optimizes a variational lower bound on mutual information using reinforcement learning. In this case, $DADS$ replaces the skill channel $p(s_{n}|s_{0},z)$ distribution with the learned variational distribution $q_{\phi}(s_{n}|s_{0},z)$, which is trained with maximum likelihood and used as part of the reward function when training the skill-conditioned policy. However, this approach will still face the same issue as those that use the learned variational distribution $q_{\phi}(z|s_{0},s_{n})$. During times when the skills are not differentiated, the reward function will not provide a strong incentive for the skills to differentiate, in turn limiting the space of skills that is learned.
@@
Additional approaches in which empowerment is used to learn skills include SNN4HRL [Florensa et al., 2017], which augments a task reward with a mutual information reward that encourages more diverse skills. LSD [Park et al., 2022] optimizes a modified version of the skill channel mutual information to force the agent to learn more dynamic and long horizon skills. However, general purpose agents need to learn skills the that are both static and dynamic as well as short and long horizon. The empowerment objective is capable of learning all of these skill types so it may offer a more promising objective.
Automated Curriculum Learning Similar to our framework are several methods that implement automated curricula of goal-conditioned RL problems. Maximum entropy-based curriculum learning methods [Pong et al., 2019, Pitis et al., 2020, Campos et al., 2020] separately optimize the two entropy terms that make up the mutual information of the skill channel. That is, they try to learn a goal space with large entropy $H(Z|s_{0})$ while also learning a goal-conditioned policy to reduce the conditional entropy $H(Z|s_{0},S_{n})$. In Skew-Fit [Pong et al., 2019], the distribution of goals tested at each phase of the curriculum is determined by a generative model, which is trained to model the distribution of states visited by the agent’s current set of skills. The generative model is skewed to be closer to a uniform distribution so that states that are new to the agent are tested more frequently. In MEGA [Pitis et al., 2020], the set of goal states tested is determined using a learned density model of the states that have been visited previously. MEGA selects states with low probabilities per the density function so the states along the frontier of the visited state space can be tested more often. In both Skew-Fit and MEGA, a goal-conditioned policy is simultaneously but separately trained to minimize $H(Z|s_{0},s_{n})$. Similarly, EDL [Campos et al., 2020] uses a three stage process to separately optimize the entropies $H(S_{n}|s_{0})$ and $H(S_{n}|s_{0},Z)$. In the exploration stage, an exploration algorithm [Lee et al., 2019] is used to try to learn a uniform distribution over the state space. In the discover phase, the state space is encoded into a latent skill space in order to try to produce a high entropy skill space. In the learning phase, a goal-conditioned policy is trained to reach goal latent states.
Another class of automated curriculum learning methods implements curricula of goal-conditioned RL problems based on learning progress. Goal GAN [Florensa et al., 2018] trains a GAN [Goodfellow et al., 2014] to output goals in which the agent has made an intermediate level of progress towards achieving. CURIOUS [Colas et al., 2019] and SAGG-RIAC [Baranes and Oudeyer, 2013] output goals in which the agent has shown improvement.
@@
Bonus-based Exploration. Empowerment is related to another class of intrinsic motivation algorithms, Bonus-based Exploration. These approaches incentivize agents to explore by augmenting the task reward with a bonus based on how novel a state is. The bonuses these methods employ include count-based bonuses [Bellemare et al., 2016, Ostrovski et al., 2017, Lobel et al., 2023], which estimate the number of times a state has been visited and provide a reward inversely proportional to this number. Another popular type of bonus is model-prediction error [Schmidhuber, 1991, Pathak et al., 2017, Burda et al., 2019] which reward states in which there are errors in the forward transition or state encoding models. One issue with bonus-based exploration methods is that it is unclear how to convert the agent’s exploration into a large space of reusable skills. A second issue, particularly with the model-prediction error bonuses, is that these methods can be difficult to implement as it is unclear how frequently to update the model that determines the bonus relative to the exploration policy.
@@ 6 Conclusion
Figure 4 shows the phase 2 results for the 3 vs. 4 level agents (i.e., 2 vs. 3 skill levels) in the 800x800 Ant domain. 800x800 represents the dimensions of the phase 2 goal space, from which goals are uniformly sampled at the start of each phase 2 episode. As in the results showed earlier in Figure 3, the y-axis shows the average minimum distance to goal given a specific period of phase 2 training. The average is over (i) four phase 1 agents that were trained and (ii) 400 goals sampled from the phase 2 goal space for each phase 1 agent. A video of a trained four level agent in the 800x800 domain is available at the following url: https://www.youtube.com/watch?v=2SW8mX-FXnc.
@@ Appendix B Comparison of Surface Area Coverage
Figure 5 compares the surface area coverage of the skills learned by Hierarchical Empowerment in the 800x800 domain to the surface area coverage in the published results of DIAYN [Eysenbach et al., 2019] and DADS [Sharma et al., 2020]. The 800x800 domain is 40,000x and $\approx 700x$ larger than the surface area covered by the skills of DIAYN (4x4) and DADS (30x30), respectively. Note that the 30x30 surface area in the DADS results was only partially covered.
@@ Appendix D Algorithm and Implementation Details
Another difference is that we optimize a regularized version of the goal-conditioned policy objective in Equation 11. Specifically, we optimize the objective
$\mathrm{E}_{z\sim p_{\phi}(z|s_{0}),s\sim\rho_{\theta}(s|s_{0},z)}[\log\mathcal{N}(s_{0}+z;s,\sigma_{0})],$ (12)
in which $\rho_{\theta}(s|s_{0},z)$ is the (improper) state visitation distribution: $\rho_{\theta}(s|s_{0},z)=\sum_{t=0}^{n-1}p(s_{0}\rightarrow s,t,z)$, in which is the $p(s_{0}\rightarrow s,t,z)$ is the probability of moving from the initial state $s_{0}$ to state $s$ in $t$ actions when the goal-conditioned policy is pursuing goal $z$. The main difference between the regularized version of the objective is that the reward $r(s_{t},z,a_{t})=\mathrm{E}_{s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})}[\log\mathcal{N}(s_{0}+z|s_{t+1},\sigma_{0})]$ occurs at each of the $n$ steps instead of just the last step. We found this to produce more stable goal-conditioned policies.
Algorithm 1 provides pseudocode for Hierarchical Empowerment. The approach is implemented by repeatedly iterating through the $k$ levels of the agent and updating the goal-conditioned actor-critic and goal space actor-critic.
Algorithm 2 provides the goal-conditioned actor-critic update function. The purpose of this function is to train the goal-conditioned policy to be better at achieving goals within and nearby the current goal space. We assume a deterministic goal-conditioned policy, so the update function uses Deterministic Policy Gradient (DPG) [Silver et al., 2014] to update the actor.
@@
     Compute reward $r_{t+1}(s_{t+1},z)=\log\mathcal{N}(s_{0}+z;s_{t+1},\sigma_{0})$
@@
Compute reward $r=\log\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})-\log p(z|a)$
@@
Appendix E Key Hyperparameters
@@
Appendix F DIAYN Visualization

### v2

@@ 4.2 Results
We present our results in a few ways. Figure 3 provides charts for the phase two performances for most of the experiments. Results of the three versus four level Ant experiment is provided in section A of the Appendix. We also provide a video presentation showing sample episodes of the trained agents at the following url: https://www.youtube.com/watch?v=OOjtC30VjPQ. In addition, we provide before and after images of the learned goal spaces in section C of the Appendix. Section K of the Appendix discusses some of the ablation experiments that were run to examine the limitations of the framework.
The experimental results support both contributions of the framework. Goal-Conditioned Empowerment substantially outperforms DIAYN, HIDIO, and DADS in the Ant Field – Small and Half Cheetah – Small domains. None of the baselines were able to solve either of the tasks. As we describe in section I of the Appendix, in our implementation of DIAYN we tracked the scale of the learned variational distribution $q_{\phi}(z|s_{0},s_{t})$. As we hypothesized, the standard deviation of the variational distribution did not decrease over time, indicating that the the skills were not specializing and targeting specific states. On the other hand, per the performance charts and video, Goal-Conditioned Empowerment was able to learn skills to achieve goals in the phase two goal space.
The experimental results also support our claim that hierarchy is helpful for computing long-term empowerment. In both the two versus three level agents (i.e., one skill level vs. two skill level agents) and in the three versus four level comparison, the agents with the additional level outperformed, often significantly, the agent with fewer levels. In addition, the surface area of the skills learned by the four level Hierarchical Empowerment agents is significantly larger than has been reported in prior empowerment-based skill-learning papers. In our largest Ant domain solved by the four level agent, the phase 2 goal space with size 800x800 is over four orders of magnitude larger than the surface area reported in the DIAYN Ant task (4x4), and over 2 orders of magnitude larger than the area (30x30) partially achieved by Dynamics-Aware Unsupervised Discovery of Skills (DADS) (Sharma et al., 2020). Section B of the Appendix provides a visual of the differences in surface area coverage among Hierarchical Empowerment, DADS, and DIAYN. In terms of wall clock time, the longest goals in the 800x800 domain required over 13 minutes of movement by a reasonably efficient policy.
@@ 5 Related Work
There are several categories of skill-learning closely related to our proposed framework (see section L of the Appendix for additional related work detail).
@@
Other Empowerment There are some additional approaches that use empowerment or modified versions of empowerment to learn skills. These include SNN4HRL (Florensa et al., 2017), which augments a task reward with a mutual information reward to encourage more diverse skills. In addition, LSD (Park et al., 2022) optimizes a modified version of the skill channel mutual information to force the agent to learn more dynamic and long horizon skills.
Automated Curriculum Learning Similar to our framework are several methods that implement automated curricula of goal-conditioned RL problems. Maximum entropy-based curriculum learning methods (Pong et al., 2019; Pitis et al., 2020; Campos et al., 2020) separately optimize the two entropy terms that make up the mutual information of the skill channel. They generally try to learn a highly entropic generative model of the states visited and use the generative model as the goal space distribution (i.e., maximize $H(Z|s_{0})$), while also separately training a goal-conditioned policy (i.e., minimize $H(Z|s_{0},S_{n})$). In addition, there are other automated curriculum learning methods that implement curricula of goal-conditioned RL problems based on learning progress (Florensa et al., 2018; Colas et al., 2019; Baranes & Oudeyer, 2013).
@@
Bonus-based Exploration. Empowerment is related to another class of intrinsic motivation algorithms, Bonus-based Exploration. These approaches incentivize agents to explore by augmenting the task reward with a bonus based on how novel a state is. The bonuses these methods employ include count-based bonuses (Bellemare et al., 2016; Ostrovski et al., 2017; Lobel et al., 2023), which estimate the number of times a state has been visited and provide a reward inversely proportional to this number. Another popular type of bonus is model-prediction error (Schmidhuber, 1991; Pathak et al., 2017; Burda et al., 2019) which reward states in which there are errors in the forward transition or state encoding models. One key issue with bonus-based exploration methods is that it is unclear how to convert the agent’s exploration into a large space of reusable skills.
@@ 6 Conclusion
Figure 4 shows the phase 2 results for the 3 vs. 4 level agents (i.e., 2 vs. 3 skill levels) in the 800x800 Ant domain. 800x800 represents the dimensions of the phase 2 goal space, from which goals are uniformly sampled at the start of each phase 2 episode. As in the results showed earlier in Figure 3, the y-axis shows the average minimum distance to goal given a specific period of phase 2 training. The average is over (i) four phase 1 agents that were trained and (ii) 400 goals sampled from the phase 2 goal space for each phase 1 agent. A video of a trained four level agent in the 800x800 domain is available at the following url: https://www.youtube.com/watch?v=Ghag1kvMgkw.
@@ Appendix B Comparison of Surface Area Coverage
Figure 5 compares the surface area coverage of the skills learned by Hierarchical Empowerment in the 800x800 domain to the surface area coverage in the published results of DIAYN (Eysenbach et al., 2019) and DADS (Sharma et al., 2020). The 800x800 domain is 40,000x and $\approx 700x$ larger than the surface area covered by the skills of DIAYN (4x4) and DADS (30x30), respectively. Note that the 30x30 surface area in the DADS results was only partially covered.
@@ Appendix D Algorithm and Implementation Details
Another difference is that we optimize a regularized version of the goal-conditioned policy objective. Specifically, we optimize the objective
$\mathbb{E}_{z\sim p_{\phi}(z|s_{0}),s\sim\rho_{\theta}(s|s_{0},z)}[\log\mathcal{N}(h(s_{0})+z;s,\sigma_{0})],$ (9)
in which $\rho_{\theta}(s|s_{0},z)$ is the (improper) state visitation distribution: $\rho_{\theta}(s|s_{0},z)=\sum_{t=0}^{n-1}p(s_{0}\rightarrow s,t,z)$, in which is the $p(s_{0}\rightarrow s,t,z)$ is the probability of moving from the initial state $s_{0}$ to state $s$ in $t$ actions when the goal-conditioned policy is pursuing goal $z$. The main difference between the regularized version of the objective is that the reward $r(s_{t},z,a_{t})=\mathbb{E}_{s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})}[\log\mathcal{N}(h(s_{0})+z|s_{t+1},\sigma_{0})]$ occurs at each of the $n$ steps instead of just the last step. We found this to produce more stable goal-conditioned policies.
Algorithm 1 provides pseudocode for Hierarchical Empowerment. The approach is implemented by repeatedly iterating through the $k$ levels of the agent and updating the goal-conditioned actor-critic and goal-space actor-critic.
Algorithm 2 provides the goal-conditioned actor-critic update function. The purpose of this function is to train the goal-conditioned policy to be better at achieving goals within and nearby the current goal space. We assume a deterministic goal-conditioned policy, so the update function uses Deterministic Policy Gradient (DPG) (Silver et al., 2014) to update the actor.
@@
     Compute reward $r_{t+1}(s_{t+1},z)=\log\mathcal{N}(h(s_{0})+z;s_{t+1},\sigma_{0})$
@@
Compute reward $r=\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0})-\log p(z|a)$
@@
Appendix E Hierarchical Architecture Detail
Figure 11: Policy architectures for level $k$ of the hierarchical agent.
Figure 11 provides some additional detail on the architectures of the goal-conditioned and goal-space policies. The figure on the left side shows the architecture for the goal-conditioned policy at level $k$, $\pi_{\theta_{k}}$. $\pi_{\theta_{k}}$ receives as input a state $s$ and goal $z$. The input is passed through the goal-conditioned policy neural network $f_{\theta_{k}}(s,z)$. $f_{\theta_{k}}(s,z)$ is then passed through a tanh activation function, which bounds $f_{\theta_{k}}(s,z)$ between $[-1,1]$. The tanh output is then multiplied by $\text{bounds}_{k-1}$, which represents the halfwidths of the level $k-1$ goal space. The goal space shift, $\text{shifts}_{k-1}$, is then added to the resulting output. $[\text{bounds}_{k-1},\text{shifts}_{k-1}]$ can be sampled from the goal-space policy $\mu_{\phi_{k-1}}$ at level $k-1$. Note that for the bottom level policy that outputs primitive actions in the range $[-1,1]$, $\text{bounds}_{-1}=1$, and $\text{shifts}_{-1}=0$.
The right side of the figure shows the architecture for a level $k$ goal-space policy. The goal-space policy $\mu_{\phi_{k}}$ at level $k$ takes as input a state $s$ and outputs the bounds (i.e., the halfwidths) and shifts of each dimension of the goal space. Note that in phase 2 of our experiments when an additional level is added to the agent in order to learn a policy over the skills, this added level only includes a goal-conditioned policy and not a goal-space policy.
Appendix F Goal Space Critic Objective
The objective for training the goal-space critic is
$J(\omega)$ $=\mathbb{E}_{(s_{0},\epsilon,a)\sim\beta}[(R_{\omega}(s_{0},\epsilon,a)-\text{Target})^{2}],$ (10)
Target $=\mathbb{E}_{s_{n}\sim p(s_{n}|s_{0},z)}[\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0})-\log p(z|s_{0})].$
In Equation 10, the learned critic is trained to be closer to a target value using $(s_{0},\epsilon,a)$ tuples from a buffer $\beta$, in which $s_{0}$ is a skill start state, $\epsilon$ is a unit uniform random variable from the reparameterization trick, $a$ is a goal space action (i.e., the vector of location-scale parameters for the uniform distribution goal space). The target is obtained by first executing the skill $z=g(\epsilon,a)$ and then sampling the terminal state $s_{n}$. The target is then the sum of (i) the $\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0})$, which measures how effective the goal-conditioned policy is at achieving the goal state $h(s_{0})+z$ and (ii) $-\log p(z|s_{0})$, which comes from the entropy reward that was folded in.
Appendix G Key Hyperparameters
@@
Appendix H DADS Objective
DADS optimizes a variational lower bound to a related objective to Empowerment: $I(Z;S_{1},S_{2},\dots,S_{n}|s_{0})$ (i.e., each skill should target a trajectory of states). The authors observe that a variational lower bound on this objective is equivalent to a reinforcement learning problem in which the reward is $r(s_{t},a_{t},s_{t+1})=\log\frac{q_{\phi}(s_{t+1}|s_{0},s_{t},z)}{p(s_{t+1}|s_{0},s_{t})}$, in which $q_{\phi}(s_{t+1}|s_{0},s_{t},z)$ is a learned variational distribution that seeks to learn the distribution of next states and the marginal $p(s_{t+1}|s_{0},s_{t})$ is estimated with the average $(\sum_{i=1}^{L}q_{\phi}(s_{t+1}|s_{0},s_{t},z_{i}))/L$ for $z_{i}\sim p(z).$ This reward will be higher if skills $z$ target unique regions of the state trajectory space as the variational distribution $q_{\phi}(s_{t+1}|s_{0},s_{t},z)$ will be high and the marginal probability $p(s_{t+1}|s_{0},s_{t})$ will be low. The problem again with this strategy is that in times when the skills overlap $p(s_{t+1}|s_{0},s_{t})$ will generally be high resulting in rewards that are generally low, which in turn produces little signal for the skills to differentiate.
Appendix I DIAYN Visualization
@@
Appendix J Brax Half Cheetah Gait
At the time of writing, the default forward gait from the Brax (Freeman et al., 2021a) version of Half Cheetah is noticeably different than the typical MuJoCo (Todorov et al., 2012) gait. Instead of the half cheetah using its feet to run, the half cheetah will propel itself forward by pushing off it back knee. We provide a video of the default forward gait at the following link: https://www.youtube.com/watch?v=4IKUf19GONw. Our implementation of half cheetah exacerbates this issue as we increase the number of physics steps per action so that the agent can move a reasonable amount in a smaller number $N$ actions.
Appendix K Ablation Experiments
In addition to the domains described in section 4.2 that were designed to support the goal space growth of Hierarchical Empowerment agents, we also implemented some ablation experiments that confirmed the current limitations of the framework. These includes various types of mazes. For instance, in one experiment we implemented a maze in the shape of an “H” with a horizontal hallway in the middle and two vertical hallways at each end of the horizontal hallway. The episodes start with the agent in the center of the horizontal hallway. In this experiment and in the other mazes we implemented, the agent was not able to learn skills that extend beyond the initial horizontal hallway. The likely reason is that in order to further extend the goal space to the vertical hallways, the goal space would need to include unreachable states that are north and south of the initial horizontal hallway. Although the entropy component of the goal-space reward ($-\log p(z|s_{0})$)is larger when parts of the vertical hallways are included because the goal space is larger, the other component of the reward that measures how well goals within the goal space are achieved $(\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0}))$ is lower because $s_{n}$ is further from the goal $h(s_{0})+z$.
Appendix L Related Work Detail
There is an extensive history of skill-learning approaches (Parr & Russell, 1997; Sutton et al., 1999b; Dietterich, 1999; Stolle & Precup, 2002; Simsek et al., 2005; Konidaris & Barto, 2009; Bacon et al., 2017). Here we provide some additional details not mentioned in the body of the paper due to space constraints.
Automated Curriculum Learning Similar to our framework are several methods that implement automated curricula of goal-conditioned RL problems. Maximum entropy-based curriculum learning methods (Pong et al., 2019; Pitis et al., 2020; Campos et al., 2020) separately optimize the two entropy terms that make up the mutual information of the skill channel. That is, they try to learn a goal space with large entropy $H(Z|s_{0})$ while also learning a goal-conditioned policy to reduce the conditional entropy $H(Z|s_{0},S_{n})$. In Skew-Fit (Pong et al., 2019), the distribution of goals tested at each phase of the curriculum is determined by a generative model, which is trained to model the distribution of states visited by the agent’s current set of skills. The generative model is skewed to be closer to a uniform distribution so that states that are new to the agent are tested more frequently. In MEGA (Pitis et al., 2020), the set of goal states tested is determined using a learned density model of the states that have been visited previously. MEGA selects states with low probabilities per the density function so the states along the frontier of the visited state space can be tested more often. In both Skew-Fit and MEGA, a goal-conditioned policy is simultaneously but separately trained to minimize $H(Z|s_{0},s_{n})$. Similarly, EDL (Campos et al., 2020) uses a three stage process to separately optimize the entropies $H(S_{n}|s_{0})$ and $H(S_{n}|s_{0},Z)$. In the exploration stage, an exploration algorithm (Lee et al., 2019) is used to try to learn a uniform distribution over the state space. In the discover phase, the state space is encoded into a latent skill space in order to try to produce a high entropy skill space. In the learning phase, a goal-conditioned policy is trained to reach goal latent states.
Another class of automated curriculum learning methods implements curricula of goal-conditioned RL problems based on learning progress. Goal GAN (Florensa et al., 2018) trains a GAN (Goodfellow et al., 2014) to output goals in which the agent has made an intermediate level of progress towards achieving. CURIOUS (Colas et al., 2019) and SAGG-RIAC (Baranes & Oudeyer, 2013) output goals in which the agent has shown improvement.
Bonus-based Exploration. Empowerment is related to another class of intrinsic motivation algorithms, Bonus-based Exploration. These approaches incentivize agents to explore by augmenting the task reward with a bonus based on how novel a state is. The bonuses these methods employ include count-based bonuses (Bellemare et al., 2016; Ostrovski et al., 2017; Lobel et al., 2023), which estimate the number of times a state has been visited and provide a reward inversely proportional to this number. Another popular type of bonus is model-prediction error (Schmidhuber, 1991; Pathak et al., 2017; Burda et al., 2019) which reward states in which there are errors in the forward transition or state encoding models. One issue with bonus-based exploration methods is that it is unclear how to convert the agent’s exploration into a large space of reusable skills. A second issue, particularly with the model-prediction error bonuses, is that these methods can be difficult to implement as it is unclear how frequently to update the model that determines the bonus relative to the exploration policy.